OpenStudy (anonymous):

Here are the first four terms of an arithmetic sequence. 10 16 22 28 (a) Find the 10th term of this sequence. .............................................................. (1) (b) Find an expression, in terms of n, for the nth term of the sequence.

OpenStudy (anonymous):

\[s _{n}=10+6(n-1)\]where\[n=1,2,3...\]So,\[s _{10}=10+6(10-1)=10+6*9=64\]